### v1

# EDITOR: Effective and Interpretable Prompt Inversion for Text-to-Image Diffusion Models

Abstract
Text-to-image generation models (e.g., Stable Diffusion) have achieved significant advancements, enabling the creation of high-quality and realistic images based on textual descriptions. Prompt inversion, the task of identifying the textual prompt used to generate a specific artifact, holds significant potential for applications including data attribution, model provenance, and watermarking validation. Recent studies introduced a delayed projection scheme to optimize for prompts representative of the vocabulary space, though challenges in semantic fluency and efficiency remain. Advanced image captioning models or visual large language models can generate highly interpretable prompts, but they often lack in image similarity. In this paper, we propose a prompt inversion technique called EDITOR for text-to-image diffusion models, which includes initializing embeddings using a pre-trained image captioning model, refining them through reverse-engineering in the latent space, and converting them to texts using an embedding-to-text model. Our experiments on the widely-used datasets, such as MS COCO, LAION, and Flickr, show that our method outperforms existing methods in terms of image similarity, textual alignment, prompt interpretability and generalizability. We further illustrate the application of our generated prompts in tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
1 Introduction
Figure 1: Original prompts and generated images along with inverted prompts.
Text-to-image generation models have made remarkable progress, greatly enhancing the creation of visual content. Advanced deep learning techniques enable models like Parti [1], DALL-E [2], and Stable Diffusion [3] to produce high-quality, realistic, and creative images based on textual descriptions. A key factor in successfully generating high-quality images is the design of prompts. Well-crafted prompts guide these models to generate images that match the desired details and styles. As such, prompts used to generate high-quality images have become a critical component of the ecosystem and valuable IP assets [4, 5, 6].
Prompt inversion is the task of reconstructing the textual prompt used to generate a specific image. It has various applications in trustworthy artificial intelligence (AI), such as data attribution, model provenance, and watermarking validation [7, 8, 9, 10, 6]. However, as summarized in Table 1, current methods face two main challenges: they either lack prompt interpretability or fall short in image similarity.
Table 1: Challenges of current prompt inversion.
Methods PEZ PH2P Image Captioning EDITOR
Prompt Interpretability?
Image Similarity?
C1. Prompt Interpretability? Existing prompt inversion methods [11, 12] are localized in the latent space and optimize for prompts that are representative of the vocabulary space. Specifically, As shown in 2(a), they optimize the latent embeddings without processing them through transformer layers and project them back to the nearest neighbors in the vocabulary space. As a result, the inverted prompts often lack semantic fluency and these methods are computationally inefficient, searching within discrete spaces. Prompts that are difficult to interpret also severely hinder their generalizability, greatly reducing their value.
C2. Image Similarity? A possible solution to get interpretable prompts involves directly employing advanced image captioning models or visual large language models to obtain accurate and fluent prompts. For example, BLIP-2 [13], LLaVA-1.5 [14] and Janus-Pro [15]. However, this direct generation often lacks similarity in image generation, especially when the goal is to produce a description that is maximally aligned with a specific image using a particular difuusion model.
(a) Optimization of embedding in PEZ and PH2P.
(b) Optimization of embedding in EDITOR.
Figure 2: Existing works optimize embedding before transformer layer and project them to nearest token embeddings, often producing incoherent results; EDITOR optimizes embedding after transformer layer and converts the final optimized embeddings into prompts. The red arrow indicates the backward pass in gradient-based optimization.
In contrast, we propose a prompt inversion method, called EDITOR, for text-to-image diffusion models to effectively generate prompts that are both accurate and elegant. The overall process starts by initializing the latent embedding using a pre-trained image captioning model. Then, we optimize the prompts to recover a more accurate and contextually appropriate prompt. In particular, as shown in 2(b), we first initialize the latent text embedding, process it through transformer layers to obtain the contextualized embedding, then apply gradient-based optimization to obtain the reconstructed embedding, and finally invert it back to the text space using a trained embedding-to-text model [16]. These three steps work together to ensure that the inverted prompt maintains coherence and interpretability. Figure 1 illustrates examples of our prompt inversion result and comparison with existing work [12, 11]. Compared to baselines, our method produces much more coherent and interpretable prompts. Furthermore, the reverse-engineering step ensures that the generated prompt remains semantically aligned with the target image, offering improved image quality compared to directly using an image captioning model.
Our results on the widely-used datasets (e.g., MS COCO [17], LAION [18], and Flickr [19]) under state-of-the-art text-to-image generation model (e.g., Stable Diffusion [3]) show that EDITOR is more effective in both image similarity and prompt quality than existing methods. Our contributions are summarized as follows:
  • We introduce a novel technique called EDITOR for inverting prompts in text-to-image diffusion models. Unlike existing methods that use discrete optimization with projection on vocabulary, EDITOR improves optimization effectiveness and efficiency through proper initialization and reverse-engineering in a continuous latent space.
  • The EDITOR employs a novel three-step pipeline including initialization, revere-engineering and embedding inversion. This framwork surpasses existing prompt inversion methods like PEZ [12] and PH2P [11] in prompt interpretability and outperforms image captioning models in image similarity.
  • We demonstrate that EDITOR produces semantically meaningful prompts excelling in image similarity, text alignment, and interpretability across three standard datasets. Moreover, they enable flexible downstream tasks such as cross-concept image synthesis, concept manipulation, evolutionary multi-concept generation and unsupervised segmentation.
2 Related Work
Text-to-Image Generation. Early approaches to text-to-image generation were dominated by generative adversarial networks (GANs) [20, 21] to generate style-controllable images. Inspired by the success of autoregressive transformer [22], recent methods [23, 24] generate images sequentially by modeling the conditional probability of pixels or patches.
Significant progress in high-resolution image synthesis has been made with diffusion models [25] that gradually refines an initial noisy input into a coherent image. Conditional diffusion models [2, 26] leverage text encoders [27, 28] to map texts into latent embeddings for controllable generation. To improve efficiency, latent diffusion models (LDMs) [3, 29, 30], perform the diffusion process in a compressed latent space instead of the pixel space. Given an image $x$, an encoder $\mathcal{E}(\cdot)$ compresses the image into a latent representation $z$, and a decoder $\mathcal{D}(\cdot)$ reconstructs the image back into pixel space, such that $x\approx\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$. Here, $z$ represents the latent representations of the image. During the diffusion process, noise is incrementally added to $z$, resulting in a sequence of latent representations $z_{t}$ for each timestep $t\in\{0,1,\dots,T\}$, where $T$ is the total number of timesteps. A U-Net [31] $\epsilon_{\theta}(z_{t},t)$ is trained as a denoising module to predict the noise component within $z_{t}$. In conditional text-to-image diffusion models, the denoising process is guided by conditional text inputs. A transformer-based text encoder $\mathcal{T}(\cdot)$ is employed to convert a given prompt $p$ into a corresponding latent text embedding $c:=\mathcal{T}(p)$. The conditional U-Net $\epsilon_{\theta}$ takes $(z_{t},t,c)$ as input, and the training objective of the conditional latent diffusion model is formalized as
$\mathbb{E}_{z,c,\epsilon,t}\big{[}\|\epsilon_{\theta}(z_{t},t,c)-\epsilon\|_{2}^{2}\big{]}.$ (1)
with $\epsilon_{\theta}$ predicting the noise and $\epsilon$ denoting the ground-truth noise. During inference in the conditional diffusion model, we initialize the process with a latent seed, which is used to generate random latent noise $n$ of the same shape as $z_{t}$, serving as the starting point for denoising. The process of synthesizing an image can be expressed as $x=\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(c,n))=\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(\mathcal{T}(p),n))$. where $\mathcal{R_{\epsilon_{\theta}}}$ is the generation process with $\epsilon_{\theta}$ in the latent space.
Inversion for Diffusion Models. Previous inversion techniques in diffusion models aim to enhance controllability and personalization in text-to-image generation by embedding visual concepts into the model’s textual or latent space [32, 33, 34, 35, 36]. DreamBooth [34] fine-tunes diffusion models using specific reference images to learn new visual features for personalized image generation. Textual inversion [32] encodes visual concepts as unique tokens in the model’s vocabulary, enabling customized image generation through textual prompts. [36] presents an encoder-based approach for fast personalization of text-to-image models. Similarly, ELITE [35] extends textual inversion by encoding multiple visual concepts into the textual embedding space, allowing the combination of various concepts within a single prompt. Moreover, [33] extends prompt conditioning, where different embeddings are injected into different layers of the U-net. Although these methods enhance personalization and controllability in text-to-image generation, they often lack interpretability and generalization, limiting their applicability to diverse contexts. One solution is to invert image features into natural language representations. Inspired by discrete optimization, recent works focus on optimizing textual prompts to better align with image features [12, 11][12] introduces a gradient-based discrete optimization method for hard prompt discovery with CLIP [37] and [11] proposes hard prompt optimization that aligns prompts with image features directly in diffusion models. These approaches bridge the gap between visual and textual representations, allowing for more intuitive manipulation of generated images through natural language. However, these inverted prompts often lack semantic fluency and naturalness, resulting in outputs unintelligible to humans. In contrast, our method ensures image similarity while making prompts more human-understandable, thereby enhancing their interpretability and generalizability.
Figure 3: Overview of EDITOR. Our approach comprises three steps: ① initializing the latent embedding via a pre-trained image captioning model; ② refining the embedding through reverse-engineering; ③ mapping the refined embedding to text with an embedding-to-text model. This pipeline yields coherent prompts and streamlines optimization.
3 Methodology
In this section, we present the method developed for prompt inversion in text-to-image diffusion models. We begin by outlining the problem formulation for the problem under study, followed by an introduction to our prompt inversion method.
3.1 Problem Formulation
We aim to invert prompts from text-to-image generation models. Following existing works [12, 11], we formulate the problem as follows:
Engineer’s Goal. The goal of prompt inversion for a given image $x$ is to find the prompt $p^{*}$ that generates an image $\mathcal{M}(p^{*})$ as close as possible to $x$. Formally, the objective can be expressed as constructing an inversion algorithm $\mathcal{A}:(\mathcal{M},x)\mapsto P$, where $P$ represents the text space.
Engineer’s Capability. Same as [11], we assume that we have white-box access to a latent diffusion model $\mathcal{M}$, allowing them to access intermediate outputs and compute the model’s gradients. In addition, the attacker has access to other public models and tools, which aids in more efficient prompt optimization.
3.2 Overview
Algorithm 1 Prompt Inversion for Text-to-Image Diffusion Model
Input:   Text-to-Image Diffusion Model: $\mathcal{M}$, Image Captioning Model: $\mathcal{M}_{c}$, Embedding-to-Text Model: $\mathcal{M}_{t}$, Target Image: $\bm{x}$.
Output:   Inverted Prompt
1:function Inversion($\mathcal{M},\bm{x}$)
2:     $\mathcal{T}=\mathcal{M}$.Text Encoder
3:     $\triangleright$ Initialization
4:     $p\leftarrow$ $\mathcal{M}_{c}$($\bm{x}$)
5:     $c\leftarrow$ $\mathcal{T}$($p$)
6:     $\triangleright$ Reverse-engineering
7:     for $e\leq\rm{max\_epoch}$ do
8:         $loss=\mathcal{L}\left(\mathcal{M}(\bm{c}),\bm{x}\right)$
9:         $\Delta_{\bm{c}}=\frac{\partial loss}{\partial\bm{c}}$
10:         $\bm{c}=\bm{c}-lr\cdot\Delta_{\bm{c}}$
11:     end for
12:     $\triangleright$ Embedding Inversion
13:     return $\mathcal{M}_{t}$($\bm{c}$)
14:end function
The overview of EDITOR is shown in Figure 3, consisting of three main steps: initialization, reverse-engineering, and embedding inversion. Firstly, we initialize the latent text embedding using an image captioning model, which provides a starting point with a semantically meaningful prompt, addressing the issue of isolated and incoherent tokens. Secondly, we apply reverse-engineering to this initial embedding to obtain a reconstructed embedding. Finally, we convert the reconstructed embedding back to text using an embedding-to-text model.
Algorithm. The overall gradient-based optimization algorithm is shown in Algorithm 1. We begin by using an image captioning model to transform the target image into a textual description, which is then embedded by the text encoder of the diffusion model. Next, we iteratively refine this embedding by calculating the gradient of a reconstruction loss that compares the diffusion model’s generated output against the original image, and we update the embedding accordingly over multiple epochs. Finally, we feed the refined embedding to a text-to-image model, thereby recovering an "inverted prompt" that closely aligns with the visual content of the target image.
3.3 Initialization for Latent Text Embedding
We introduce a novel initialization strategy that leverages pre-trained image captioning models to generate an initial prompt for the target image. This approach ensures that the inverted prompt begins with a semantically meaningful and contextually relevant prompt, thereby reducing the search space and aligning the embedding more closely with the target distribution. Specifically, for a given image $x$, we utilize an image captioning model to generate an initial prompt $p_{0}$. This initial prompt is then encoded into a starting point fot optimization using the text encoder $\mathcal{T}$, which maps the text into the latent space of the diffusion model.
It is worth noting that the initialization strategy is unique to our method, benefiting from the optimization in a continuous latent space. Unlike hard prompt optimization methods such as PEZ [12] and PH2P [11], our approach does not significantly disrupt the syntactic structure of the initial prompt. In contrast, as shown in section 6, even with initialization, PEZ and PH2P produce token combinations with high perplexity.
3.4 Reverse-engineering for Latent Text Embedding
With the initial prompt, we reconstruct latent text embeddings using input reverse-engineering. Given a target image $x$ and a latent diffusion model, the objective is to find an embedding $c^{*}$ such that the generated image $\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(c^{*},n))$ minimizes the distance to $x$. This input reverse-engineering task is formulated as:
$c^{*}=\arg\min_{c}\mathcal{L}(\mathcal{D}(\mathcal{R_{\epsilon_{\theta}}}(c,n)),x).$ (2)
where $\mathcal{L}$ is a distance metric that measures the similarity between two images, referred to as the reconstruction loss. The latent random noise $n$ is fixed, and the parameters of the latent diffusion model are held constant.
We use gradient-based optimization to iteratively search for the optimal embedding $c^{*}$ by updating the embedding based on the gradient of the reconstruction loss. For each step, the embedding $c$ is updated as follows:
[…]
where $lr$ is the learning rate. This process continues until the maximum number of epochs is reached.
In contrast to existing methods [12, 11] that optimize the embeddings before passing through the transformer layers and then project them onto the corresponding vocabulary, we directly optimize the output of the text encoder in the latent space. This approach offers two key advantages: (1) it improves optimization efficiency by bypassing the projection step, which can be computationally expensive and slow, and it ensures better semantic consistency and coherence. As shown in  Table 14, EDITOR improves optimization efficiency by more than double compared to PH2P; (2) by optimizing the text encoder’s output, which has already undergone deeper semantic processing, the generated text remains more natural, contextually relevant, and grammatically correct, while also avoiding the disjointed or redundant word choices that can arise in discrete optimization.
3.5 Embedding Inversion
It is also challenging to get text prompts from the reconstructed embedding, as the embedding is continuous and highly contextualized, making it impossible to be projected to the nearest token in the vocabulary, as is done in PEZ [12] and PH2P [11]. To solve this problem, we leverage an embedding-to-text model [16] to convert the reconstructed embedding back into text prompts. Let $c\in\mathbb{R}^{d}$ be a target (latent) embedding. We first train a zero-step model $M_{\mathrm{zero}}$ that directly maps $c$ to text $x$, estimating the conditional probability $p_{\mathrm{zero}}(x\mid c)$. This training uses a maximum likelihood objective:
$\mathcal{L}_{\text{ML}}=-\sum{(x,c)}\log p_{\mathrm{zero}}\bigl{(}x\mid c\bigr{)}.$ (4)
Once trained, $M_{\mathrm{zero}}$ generates an initial hypothesis $\hat{x}$ by sampling from $p_{\mathrm{zero}}(x\mid c)$. Next, we introduce a correction model $M_{\mathrm{corr}}$ to iteratively refine $\hat{x}$. In each refinement step $k$, the model produces a revised text $x^{(k)}$ conditioned on 1) the target embedding $c$, 2) the current hypothesis $x^{(k-1)}$, and 3) the embedding $\mathcal{T}\bigl{(}x^{(k-1)}\bigr{)}$. Formally:
$p_{\mathrm{corr}}\bigl{(}x^{(k)}\mid c;x^{(k-1)};\mathcal{T}(x^{(k-1)})\bigr{)}.$ (5)
Here $\mathcal{T}(\cdot)$ denotes the text encoder that maps text to the same latent space as $c$. The correction model is realized with a transformer-based encoder-decoder and is trained via cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\sum{(x^{(k)},c)}\log p_{\mathrm{corr}}(x^{(k)}|c).$ (6)
During inference, beam search is applied to generate multiple refined hypotheses, and the final output is selected to minimize the distance $\bigl{\|}\,\mathcal{T}(x)\,-\,c\,\bigr{\|}$ with respect to the ground-truth embedding $c$.
Compared to PEZ [12] and PH2P [11], our embedding-to-text process significantly improves the interpretability of the inverted prompt. Due to the attention mechanism, the embedding at each position can incorporate information from previous tokens, making it more likely that phrases aligned with the image’s semantics are preserved in the generated text, rather than just individual words. Additionally, when the optimized embedding introduces new or uninterpretable features, the highly interpretable training data enables the model to convert these features into more understandable and coherent text.
Table 2: Image similarity comparison with baselines.
Dataset Metric PEZ PH2P EDITOR
MS COCO CLIP Similarity $\uparrow$ 0.745 0.789 0.796
LPIPS Similarity $\downarrow$ 0.477 0.467 0.414
LAION CLIP Similarity $\uparrow$ 0.756 0.826 0.826
LPIPS Similarity $\downarrow$ 0.455 0.427 0.401
Flickr CLIP Similarity $\uparrow$ 0.722 0.755 0.776
LPIPS Similarity $\downarrow$ 0.442 0.439 0.424
Table 3: Prompt quality comparison with baselines.
Dataset Metric PEZ PH2P EDITOR
MS COCO Precision $\uparrow$ 0.791 0.801 0.900
[…]
PPL $\downarrow$ 8,837.797 11,078.994 80.659
LAION Precision $\uparrow$ 0.783 0.794 0.841
[…]
PPL $\downarrow$ 8040.401 9621.671 187.299
Flickr Precision $\uparrow$ 0.774 0.798 0.892
[…]
PPL $\downarrow$ 9,425.353 11,804.100 90.320
4 Experiment Setup
EDITOR is implemented with Python 3.10 and PyTorch 2.1.0. We conducted all experiments on a Ubuntu 20.04 server equipped with 8 A100 GPUs (one experiment/GPU).
Datasets and Models. We evaluate our method on three widely-used datasets: MS COCO [17], which contains diverse natural images with detailed annotations; LAION [18], a large-scale dataset of image-text pairs; and Flickr [19], which consists of diverse images, including everyday life scenes, natural landscapes, animals, urban settings, and human activities. For each dataset, we randomly sample 100 images for evaluation. We employ Stable Diffusion v1.5 [3] as our target generative model for prompt inversion experiments.
Evaluation Metrics. We employ multiple metrics to comprehensively evaluate the performance of prompt inversion:
  • Image Similarity: We utilize CLIP [38] and LPIPS [39] scores to measure the visual similarity between generated and target images, assessing the reconstruction quality.
  • Textual Alignment: BERTScore [40] is used to evaluate semantic consistency between inverted and ground-truth prompts, computing Precision, Recall, and F1 metrics.
  • Prompt Interpretability: We calculate Perplexity Scores (PPL) to assess the linguistic coherence of inverted prompts, where lower scores indicate more natural and comprehensible text.
Baseline Methods. We compare our approach with state-of-the-art prompt inversion methods: PEZ [12], which inverts prompts with image-text similarity matching, and PH2P [11], which employs a delayed projection scheme to optimize prompts.
5 Experiment Results
Table 4: Images generated with inverted prompts. For a given target image(left), inverted prompts from PEZ [12], PH2P [11] and our method are used to prompt Stable Diffusion (v1.5) [3], DALL-E 3 [41] and Ideogram 2.0 (right).
Target Image Generated Images Target Image Generated Images
Stable Diffusion Dall-E 3 Ideogram 2.0 Stable Diffusion Dall-E 3 Ideogram 2.0
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEZ: beach firing fortnight !), lgbti takeaways deploy acquisition labs recognised xerirrigation PEZ: hiber❄ unforgettable ophthalstowe maine mansion during giftideas indulge travelunitedstates
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PH2P: petersburg adid micropoetry beach houses shack houses spend out sunbathing living foodand PH2P: ride mice liking dorchester russian mansion attemp population ? yann reduced exception
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
EDITOR: Affraies sitting in front of gloomy and colorful shacks on a beach EDITOR: Afraed House with a lit porch and balcony in the snow
Target Image Generated Images Target Image Generated Images
Stable Diffusion Dall-E 3 Ideogram 2.0 Stable Diffusion Dall-E 3 Ideogram 2.0
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PEZ: contemplating forests reignhormone blame millennials cleanenergy !), scouts immunology scientists PEZ: ajes pocheña ados psv skull pave gold ceramic watch mal ados
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PH2P: outdoors hamp resignation sometimes backyard smokey dreaming relationships forest forest mountain exploring PH2P: tattoo watches sugar skull skull true housewives grand ( ese patriarch
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
EDITOR: a woman in a red dress standing on a fallen tree in the woods EDITOR: a sugar skull watch in a pink box
Image Similarity. We evaluate the effectiveness of the inverted prompts generated by EDITOR for image generation. For a given target image, we first perform inversion to recover the corresponding prompt and assess the relevance of the images generated by the diffusion model using the recovered prompt. We compare EDITOR with PEZ [12] and PH2P [11] under different datasets. The results are shown in subsection 3.5. For CLIP similarity, EDITOR achieves the highest scores. In detail, the average CLIP similarity score under MSCOCO, LAION and Flickr are 0.794, 0.826 and 0.776, respectively, outperforming PEZ and PH2P. Furthermore, EDITOR achieves a value of 0.414 under MSCOCO, 0.401 under LAION and 0.424 under Flickr in terms of the LPIPS similarity, consistently outperforming the baselines. The results demonstrate that EDITOR’s prompts achieve high reconstruction quality in terms of similarity to the target images.
Textual Alignment. We evaluate the semantic similarity between the inverted prompts and the ground-truth prompts using BERTScore [40]. As shown in subsection 3.5, EDITOR achieves consistently higher precision, recall, and F1 scores compared to the baseline methods across all datasets. Notably, under MS COCO, LAION and Flickr, EDITOR achieves a precision of 0.900, 0.841, and 0.892, surpassing the baselines by a value of 0.099, 0.047 and 0.094. When considering recall, EDITOR also excels across datasets. In detail, EDITOR can achieve a recall score of 0.920, 0.824 and 0.915 under all three datasets, respectively, outperforming the baselines. These results illustrate the effectiveness of our approach in achieving semantically meaningful text recovery that aligns closely with the ground-truth prompts.
Prompt Interpretability. We further assess the interpretability of the recovered prompts by computing their perplexity scores (PPL). Lower perplexity indicates that the inverted prompts are more fluent, natural, and interpretable. Results in subsection 3.5 show that our approach achieves significantly lower perplexity scores than PEZ and PH2P. Specifically, EDITOR achieves a PPL score of 80.659 under MS COCO, 187.299 under LAION, and 90.320 under Flickr, which far outperform the scores of PEZ (8,837.797 under MS COCO, 8,008.401 under LAION, 9,425.353 under Flickr) and PH2P (11,078.994 under MS COCO, 9,621.671 under LAION, 11,804.100 under Flickr). On MS COCO and Flickr, EDITOR shows 110% and 104% improvement compared to the baselines, while on LAION, it demonstrates 42% better performance. The vast gap in performance highlights the exceptional ability of EDITOR to recover prompts that are not only semantically aligned with the ground-truth but also significantly more fluent and interpretable.
Comparison with Image Captioning Models. To demonstrate the superior performance of EDITOR compared to state-of-the-art image captioning models or vision large language models including LLaVa-1.5-7B, Janus-Pro-7B, BLIP-Large, BLIP2-OPT-2.7B, and GIT, we evaluated their differences in image generation quality. As shown in Table 5, EDITOR achieves the highest CLIP Similarity score (0.776) and the lowest LPIPS Similarity score (0.424) when compared to other image captioning models like LLaVa-1.5-7B [14], Janus-Pro-7B [15], BLIP-large [42], BLIP2-OPT-2.7B [13], and GIT-large [43]. This indicates that the prompts generated by EDITOR have a stronger connection to the visual content of the images and are more consistent with human-labeled captions.
Figure 4: Comparison of CLIP-Similarity for initial and inverted prompts across Datasets.
Figure 5: Comparison of CLIP-Similarity for initial and inverted prompts across image captioning models.
Table 5: Comparison with advanced image captioning models or vLLMs.
Metric EDITOR LLaVA-1.5-7B Janus-Pro-7B BLIP-Large BLIP2-OPT-2.7B GIT
[…]
An important characteristic of EDITOR is its ability to optimize captions generated by an image captioning model in conjunction with a target diffusion model. This leads to improved text-image alignment. Figure 4 shows the enhancements achieved using EDITOR across datasets. The results clearly demonstrate that EDITOR significantly enhances the similarity score between generated images and the original images, increasing the CLIP similarity by at least 5%. Figure 5 shows the improvements achieved using EDITOR across different image captioning models. The results indicate that EDITOR can significantly improve the quality of generated images across various captioning models. What’s more, we can fully believe that if image understanding models or captioning models have nice performance, EDITOR can achieve better results based on them.
Ablation Studies. To show the importance of initialization in EDITOR, we compared random initialization and image captioning model initialization on a Flickr subset. As shown in section 6, random prompts hinder convergence, often leading to poor performance and local optima traps. Our initialization is crucial for guiding the algorithm to accurate, optimal solutions. For fairness, we also initialized PEZ and PH2P. As shown in section 6, initialization only slightly improved PEZ’s textual Alignment and prompt interpretability, with almost no impact on other aspects for both methods. This is because PEZ and PH2P are discrete optimization algorithms; even if they start with interpretable prompts, the optimization process may gradually replace them with high-perplexity token combinations. This also demonstrates that our method has greater potential than existing baselines, particularly in its compatibility with initialization. For additional ablation studies, refer to Appendix C.
Qualitative Results. We show example inverted prompts towards four target images and corresponding generations with three models: Stable Diffusion [3], DALL-E 3 [41], and Ideogram 2.0 in  Table 4. EDITOR produces more accurate, readable prompts that generate images closely aligned with the target images, demonstrating strong generalizability across models. In contrast, prompts from PEZ and PH2P tend to be less coherent, often resulting in images that do not align well with the intended concepts, highlighting the superior precision and versatility of our approach.
6 Application of the generated prompts
Figure 6: Application of cross-concept image synthesis.
Figure 7: Application of removal or replacement of objects.
Cross-Concept Image Synthesis. Our method combines multiple prompts into a unified prompt, enabling seamless image fusion. This supports applications like composite image generation and multi-concept scene creation, preserving semantic and visual coherence. For example, Figure 6 shows the fusion of two prompts: "a woman in a red dress standing on a fallen tree in the woods" and “colorful shacks on a beach". The resulting image integrates the woman into a beach shack scene, maintaining prompt accuracy and visual integrity.
Removal or Replacement of Concepts. Our prompt inversion method enables removal or replacement of image concepts. Unlike PEZ [12] and PH2P [11], which produce hard-to-interpret prompts, our method generates clear, human-readable prompts, simplifying token identification for concept editing. This improves flexibility in image manipulation. For example, Figure 7 shows removing "trees" from the prompt eliminates them, while replacing "trees" with "fence" substitutes a fence, demonstrating effective prompt-based editing. For more applications including evolutionary multi-concept generation and unsupervised segmentation, refer to Appendix B.
Table 6: Impact of initialization.
Metric EDITOR w/o initialization EDITOR
[…]
Prompt F1$\uparrow$ 0.829 0.873
[…]
Table 7: Results of PEZ and PH2P with initialization.
Metric PEZ PEZ+Initialization PH2P PH2P+Initialization
[…]
Prompt F1$\uparrow$ 0.820 0.867 0.819 0.818
[…]
7 Conclusion
In this paper, we introduced a prompt inversion approach for text-to-image diffusion models that generates high-quality, semantically aligned and interpretable prompts. Our method, validated on datasets like MS COCO, LAION, and Flickr, outperforms existing techniques in image similarity, textual alignment, and efficiency. Additionally, the generated prompts are versatile and applicable to tasks such as cross-concept image synthesis and object removal or replacement. These results demonstrate that prompt inversion is a promising technique for improving the performance and versatility of text-to-image generation models.
Appendix A Details about the E2T Model
Table 8: Effectiveness of the Embedding-to-Text (E2T) Model.
Precision Recall F1 Score
0.968 0.971 0.969
For the embedding-to-text (E2T) model, we use T5-base [44] as the backbone, and the training was performed on the MSMARCO corpus [45] with a batch size of 32, a learning rate of $1\times 10^{-3}$, a maximum token length of 32, an epoch number of 60 for the zero-step model and an epoch number of 35 for the correction model. We test the E2T model using 2,500 prompts selected from the MS COCO dataset [17], with evaluation metrics based on BertScore [40], including Precision (P), Recall (R), and F1-score. The results was shown in Table 8. Our E2T model achieves over 0.968 in Precision, Recall, and F1-score, demonstrating that our model can accurately reverse the embedding back into the original prompt with high precision.
Appendix B Applications of Prompt Inversion
Figure 8: Application of cross-concept image synthesis.
Cross-Concept Image Synthesis. Our method enables the concatenation of multiple prompts into a single, unified prompt, allowing for the seamless fusion of corresponding images. This ability enhances creative possibilities, enabling applications such as composite image generation, multi-concept scene creation, and image synthesis from diverse ideas, all while maintaining semantic integrity and visual coherence across the fused images. For example, Figure 8 demonstrates the fusion of two distinct visual concepts using concatenated prompts. The first part of the prompt describes “a woman in a red dress standing on a fallen tree in the woods" while the second part introduces “colorful shacks on a beach". By combining these two prompts, our method successfully generates an image where the woman in the red dress is seamlessly integrated into a new environment, sitting in front of colorful beach shacks. This showcases our method’s ability to merge different concepts and produce a coherent, contextually aligned image that maintains both the accuracy of the individual prompts and the overall visual integrity.
Removal or Replacement of Concepts. Our prompt inversion method can also be applied to the removal or replacement of concepts within an image. Unlike PEZ [12] and PH2P [11], which often generate prompts that are difficult to interpret, making it challenging to identify the corresponding tokens for the objects to be replaced, our method generates prompts that are fully human-readable and understandable. This clarity allows for easy identification of the specific tokens associated with the concepts, facilitating the seamless replacement or removal of those concepts within the image. This advantage enhances the flexibility and usability of our approach in image manipulation tasks. As shown in Figure 9, removing the word "egg" from the prompt eliminates them from the image, while replacing "egg" with "broccoli" substitutes the egg with a broccoli. This demonstrates the flexibility of our method in editing images via prompt modification.
Figure 9: Application of removal or replacement of objects.
Unsupervised Segmentation. EDITOR can be applied to unsupervised semantic segmentation by leveraging cross-attention maps to create segmentation masks [46, 47]. EDITOR generates prompts capturing key concepts directly from a diffusion model, without external data. For a target image, we use EDITOR to produce the inverted prompt, and as shown in 10, the resulting cross-attention maps align tokens with relevant image regions, enabling accurate segmentation.
Figure 10: Application of unsupervised segmentation.
Evolutionary Multi-concept Generation. Evolutionary multi-concept image generation is a process that enables users to create complex images by iteratively combining concepts from multiple sample images using prompt inversion. Consider a example in 11, we used EDITOR to get two prompts from two images: "a giraffe burger with bacon and cheese and french fries" and "there is a bowl of spinach and some tomatoes in it." We then combined these two prompts to create a new image, which resulted in a new prompt through EDITOR: "Burger with fries on a plate with salad and tomatoes on it.".
Figure 11: Application of evolutionary multi-concept generation.
Appendix C Ablation studies
Impact of Initial Prompt Length. Since our prompt inversion method starts with an initial prompt, we explore the impact of the initial prompt length on the inversion performance. We vary the initial prompt length from 12 to 20 tokens, and the results are shown in Appendix C. The results show that an initialization length of 16 tokens achieves the best performance. Shorter lengths (e.g., 12 tokens) tend to lack sufficient semantic information, making it harder to accurately reconstruct prompts. In contrast, longer prompts (e.g., 20 tokens) provide richer semantic information, leading to improved textual alignment and prompt interpretability. For instance, EDITOR achieves a prompt precision of 0.884 with 20 tokens, surpassing the precision of 0.852 achieves with 12 tokens. However, longer prompts introduce excessive complexity, making the optimization process more challenging. Compared with 20 tokens, the result of 16 tokens achieves a higher CLIP Similarity of 0.789 and better performance in textual alignment and prompt interpretability. At the same time, this length also offers a lower optimization complexity. As a result, an appropriate initialization length such as 16 tokens could be selected to balance semantic richness and optimization efficiency.
Table 9: Impact of initial text length.
Metric 12 tokens 16 tokens 20 tokens
[…]
Prompt F1$\uparrow$ 0.858 0.873 0.895
[…]
Table 10: Impact of interation.
Metric 500 1000 1500 2000
CLIP Similarity $\uparrow$ 0.784 0.796 0.799 0.802
[…]
Prompt F1 $\uparrow$ 0.908 0.908 0.904 0.903
Prompt Perplexity $\downarrow$ 70.364 80.659 83.457 87.907
Impact of the Number of Iteration. We explore the impact of the number of iteration during the optimization for embedding inversion in Appendix C. Generally, increasing the number of iterations improves the similarity between the generated image and the target image, which is the goal of our optimization. However, the improvement slows down over time, meaning that additional iterations result in smaller gains. Additionally, with more iterations, the quality of the inverted prompt decreases. This happens because over-optimization introduces complex or unclear tokens and semantic structures, making the prompt harder to interpret. To balance image similarity, prompt quality, and computational cost, we set the default number of iterations to 1000.
Impact of the Number of Denoising Step. EDITOR can invert the prompts for text-to-image diffusion models with small denoising steps. We explore the impact of denoising step number from 5 to 15 steps. As shown in Appendix C, the difference in performance across different step numbers is relatively small, indicating that the number of denoising steps has limited influence on the image similarity and the quality of prompts. However, using fewer steps, such as 5 steps, significantly reduces optimization complexity and computational cost, improving overall efficiency. In Stable Diffusion, denoising steps below 5 often lead to insufficient noise reduction, which impacts image quality. Therefore, we set the default number of denoising step to 5.
Table 11: Impact of denoising step.
Metric 5 steps 10 steps 15 steps
[…]
Prompt F1$\uparrow$ 0.908 0.904 0.903
[…]
Table 12: Impact of different image captioning model.
Metric GIT-large BLIP-large BLIP2-OPT-2.7B
[…]
LPIPS Similarity $\downarrow$ 0.427 0.414 0.467
[…]
Prompt F1$\uparrow$ 0.889 0.908 0.894
[…]
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Appendix C using three state-of-the-art open-source models: GIT-large [43], BLIP-large [48] and BLIP2-OPT-2.7B [42]. Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Different Image Captioning Models. To justify the image captioning model choices for the prompt initialization, we show the results in Appendix C using three state-of-the-art open-source models: GIT-large [43], BLIP-large [48] and BLIP2-OPT-2.7B [42]. Among the models, GIT-large shows the weakest performance. Notably, BLIP2-OPT-2.7b leverages a large language model, resulting in more sophisticated initialized prompts that achieve better textual alignment with the ground-truth. BLIP-large achieves better image quality with higher CLIP Similarity. We select BLIP-large as our default image captioning model. It is more lightweight compared to BLIP2-OPT-2.7b while still maintaining excellent performance.
Impact of Random Noise Initialization. The choice of random noise seed in diffusion models significantly affects the output. We further investigated the influence of seed values used for initializing random noise on the results of our method. As shown in Table 13, the impact is minimal. This demonstrates the robustness of our approach and the ability of the model to effectively manage and mitigate the effects of random noise during the inversion process.
Table 13: Impact of random noise initialization.
Metric 0 34 258 423
[…]
Prompt F1$\uparrow$ 0.894 0.908 0.893 0.900
[…]
Appendix D Efficiency
Table 14: Efficiency comparison between PH2P and EDITOR.
PH2P EDITOR
3482s 1369s
To run an inversion with 1000 iterations on 1 A100 GPU, the optimization times are shown in Table 14. EDITOR improves optimization efficiency by more than double compared to PH2P. We did not compare PEZ because it is based on CLIP inversion, whereas PH2P and EDITOR are based on Stable Diffusion.
Appendix E Broader Impacts
As the value of prompts in text-to-image generation models increases,The risk of prompt stealing, which could infringe upon the intellectual property of prompt engineers and undermine the commercial model of the prompt market [6, 10].In this paper, we propose a method for prompt inversion, which can reverse a target image into a high-quality prompt on text-to-image (T2I) generation models. We are confident that our approach can contribute to a better understanding of this emerging issue and provide solutions to protect the intellectual property of prompt engineers.
Our method also shows potential applications in data attribution and model provenance. The intuition is that if an image is generated by a specific diffusion model or originates from its training data, performing prompt inversion on that image using its source model will yield higher image similarity. By designing appropriate validation pipeline, we can use EDITOR to achieve data attribution and model provenance.
Appendix F Limitations
EDITOR produces prompts that are more interpretable and accurate than those of existing methods, and the resulting prompts can generate images more accurately than advanced image captioning models or visual large language models (vLLMs). However, more advanced diffusion models, such as Stable Diffusion 3 [49], utilize both the T5 model and a CLIP-based encoder model to achieve text-controlled generation. The EDITOR method can extend its capabilities to perform embedding optimization under multiple text encoders, enhancing its generality and performance on advanced models.
Appendix G Additional Results
Table 15: Additional examples of prompt inversion
Original Prompt Inverted Prompt
a photograph of an astronaut riding a horse an astronaut rides on a horse in space
An ancient ruins overgrown with lush greenery Ancient stone ruins standing in lush green, arid setting
A group of people taking photographs with cellphones A bunch of people are taking selfies with their cell phones
two cows sitting on some dirt and grass You can see that there are two cows standing in the grass
A slice of pie on top of a white plate near a fork there’s a piece of cake on a plate with a fork
Crochet Baby Bunny Ear Hat and Bootie Set - Light Blue crochet bunny hat and booties
Figure 12: Additional images with invert prompts